Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
bersanetti, ruggi - 12:01 Saturday 30 August 2025 (67600) Print this report
Increased longitudinal corrections & dithering signals

Since yesterday we are observing an anomalous behaviour in some corrections signals: looking at an unlock, many of them are similar to the one in Figure 1, where the ITF simply drifts away. However, a weird thing noticeable is a big difference in the corrections of WE/NE, when the former has much higher values at low frequency (mode at ~0.7 Hz?).

Looking at the spectra in Figure 2 the difference is visible; it is also visible the opposite behaviour of the MAR corrections (NI noisier than WI), although in a higher frequency range.

Comparing the spectra with the ones from the other night (after the change in behaviour around midnight), we can see the corrections being much noisier below 10 Hz, I don't know if this can be attributed to the weather conditions alone. Some of the spectra show structures similar to scattered light, and we can see in the 6-10 Hz region the dithering lines being completely covered.

Looking at DARM itself, where those signals are demodulated from, in Figure 4 there are three subsets of data of the same 2 minutes I used for the previous plots: the noise at low frequency is very non stationary, and going towards the unlock it becomes much higher, spoiling most probably the dithering signals used for the alignment of the soft modes.

This was observed also yesterday, when unrealistically high signals were observed before an unlock; already yesterday Paolo put in place a clipping for the signals used by the alignment loops, but this cannot fix the quality of the signals.

Images attached to this report
Comments to this report:
ruggi - 12:51 Saturday 30 August 2025 (67601) Print this report

I add a few plots that I was preparing for a similar entry. Fig 1 and fig 2 show the behaviour of COMMp, spoiled by the quality of the dither signals. The clip at +- 10 added yesterday on the dither signals reduced a bit the effect on the alignement, but the quality was still quite bad because probably this morning the disturbance was a bit larger than yesterday. Now the seismic conditions are getting better: we can hope to see also the alignment improving.

To explain better what is happening to dither signals: they are built demodulating the longtudinal correction on each mirror at the frequency of injected angular lines on the same mirror. It was a good strategy in the past, when the corrections were dominated by DARM (and DARM was better below 10 Hz). The control of CARM SLOW changed a bit the situation: its correction signal can be larger than DARM correction, because of sensing noise. In that case, the demodulated signal become blind to the angular lines. As we can see in fig 3, CARM SLOW is sensitive to seismic noise and wind. We should check if this sensitivity is worsened recently. I remember about fringes disturbing the RFC error signal in the past, which could be cured improving the efficency of the Faraday isolator, or RFC aligment, or INJ general alignment.

A simple solution for improving the dither signals is to increase the line amplitude: a side effect is a larger injection of noise on DARM at the duble frequency, around 15 Hz. That frequency region is useless for the BNS, so the cost would be low.

There are better solutions: the modulation/demodulation strategy can be for sure improved a lot, using a signal just based on DARM, but a bit of work is required. We also know that the centering signals generated by the cameras are good enough and could replace the dither signals.

Images attached to this comment
bersanetti - 21:38 Saturday 30 August 2025 (67603) Print this report

Starting from the lock at 18:09:40 UTC, the amplitude of the dithering lines on the TMs has been increased of a factor 2.5 (same amplitude as in LOW_NOISE_2) for Science Mode; to be verified if the change has an impact or not on the goodness of the signals.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×